A Complete Bibliography of Publications in Lisp and Symbolic Computation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.10

Title word cross-reference

$47.50 [4]. λ [93]. R^{2} [114].

-Calculus [93].

1991 [50].

2d [4, 15].

3D [89].

Abstract [86, 128]. Abstraction [62]. Acid [89]. Adding [119]. Addison [19, 25]. Addison-Wesley [19, 25]. Algebraic [17]. Algorithm [122]. Algorithms [85]. Analysis [20, 36, 46, 88, 98, 101, 102, 110]. Answer [50]. Applicable [13].

Application [88, 89]. Applications [69, 145]. Applicative [6]. Approach [30, 36, 48, 70]. Architecture [27, 52, 114]. Aren’t [18]. Arity [30]. Arrays [16]. Artificial [15]. Assignment [78, 142]. Associates [15]. Author [140]. Automatic [20, 123].

Baker [50]. Balancing [68]. Based [29, 45, 84, 108, 111, 128]. Basis [85]. Benchmark [40]. Benjamin [41]. Benjamin/Cummings [41]. better [113]. Binding [98, 101]. Binding-Time [98]. Book [4, 15, 19, 25, 33, 41]. Boston [4]. Bromley [4]. Browser [39].

Calculus [76, 93, 110]. Call [82]. Call-by-need [82]. Callee [59]. Callee-save [59]. Calls [21]. Categorical [109]. Cells [5]. Charniak [15]. Chinese [8]. CL [90]. Class [9, 38, 64]. Classes [44]. CLOS [19, 119]. Code [7]. Collection [27]. Combination [17, 39]. Combinator [32]. COMMON [10–12, 17, 19, 25, 28, 29, 33, 41, 58, 118]. Company [25, 41]. Compilation [57]. Compiler [13, 96, 143]. Compiling [37]. Composable [81]. Computation [41, 85]. Computing [114]. Concepts [11]. Concurrency [71]. Concurrent [54, 55, 86]. Conjunction [35]. Conjunction-Type [35]. CONSAT [87]. Considered [108]. Constraint [60, 87, 89]. Continuation [59, 77, 82, 84]. Continuation-Passing [59, 77, 82]. Continuations [74, 81, 127]. Control [24, 84]. Conversion [78, 142]. Corrigendum [142]. Cost [116]. Counting [124]. CPS [78, 93, 142]. Critique [49]. Cummings [41].

Data [85, 91]. David [41]. Debugger [23]. Debugging [6, 61]. Decision [58]. Defining [118]. Definition [48, 49]. Delimiters [24]. Demonstration [124]. Dependent [114]. Description [1]. Descriptions [75]. Designing [39]. Determination [89]. Dialect [2]. Dictionaries [9]. Dictionary [100]. Dictionary-Free [100]. Dijon [145]. DIN [48, 49]. Direct [21]. Discoveries [74]. Distributed [27, 54, 85, 91]. DPOS [56]. Duplication [113]. Dynamically [45, 46]. Dynamically-Typed [45, 46].

ed [4, 15]. Editor [97]. Editorial [65, 121, 129]. Education [67]. Efficient [45]. Encapsulation [43]. Endpaper [5, 28, 40]. Environment [8, 56, 61]. Environments [38]. Erlbaum [15]. Essence [99]. Eta [99]. Eta-Expansion [99]. Euclidian [122]. EULISP [66–68, 70–72]. Evaluation [7, 32, 63, 99, 100, 113, 122, 126, 143]. Evaluator [13, 26]. Execution [37].

Expansion [3, 99]. Expansion-Passing [3]. Experiences [114]. Experiment [143]. Experiments [13]. Extended [46]. Extension [26, 72]. Extents [64].

Fexprs [125]. First [38, 64]. First-Class [38, 64]. Foreword [97]. Formal [75]. France [145]. Franz [25]. Free [38, 100]. FRPOLY [40]. Function [5, 21]. Functional [14, 16, 22, 32, 57, 61]. Functions [12].

Garbage [27]. General [3, 52].

General-Purpose [52]. Generation

[13, 143]. Generator [123, 143]. Gentle [41]. Gentler [41]. Global [108]. Graph [90]. Grasper [90]. Grasper-CL [90]. Gröbner [85]. Guest [97]. Guide [4, 19].

Hall [33]. Hank [4]. hardbound [4, 15]. Hardware [124]. Haskell [104]. Headers [93]. Heap [124]. Hierarchies [24, 60]. Higher [115]. Higher-Order [115].

Imperative [105, 110]. Implementation [29, 45, 57, 84, 92, 94, 144]. Implementing [54, 118]. Index [140, 141]. Inference [70]. Inheritance [43]. Integrating [14, 31]. Intelligence [15]. Interactive [117]. Interface [11, 12]. Interpretation [57, 106, 128]. Interpretation-Compilation [57]. Interpreter [86, 112]. Interprocedural [20]. Introduction [41, 47, 51, 73, 79, 103, 108]. Introspection [111]. Issue [107]. Issues [5]. Iteration [80]. Iterative [46].

Jersey [15]. July [50].

Keene [19]. Kernel [48, 49]. Kinder [41]. Kluwer [4].

L [33]. Lambda [76, 110].

Lambda-Calculus [76]. Lamson [4].

Language [45, 57, 75, 106, 111]. Languages [6, 16, 61, 113]. Lawrence [15]. Lazy [7, 61]. Level [123]. LISP [4, 8, 10–12, 17, 19, 21, 25, 28, 29, 33, 37, 41, 48, 49, 58, 91, 118]. Local [109]. Logic [14, 31, 53, 57, 84]. Logical [26, 32, 36]. LogScheme [31]. Loop [93]. Lore [4].

Machine [4]. Macro [3]. Management [90, 116]. Mark [27]. Massachusetts [19]. May [145]. Mayfly [52–54]. McDermott [15]. Measuring [116]. Mechanism [3]. Meehan [15]. Message [46]. Metacircular [26]. Metalanguage [56]. Metaobject [68, 119]. Method [17, 39, 115]. Milner [33]. Mix [13]. ML [35, 98, 120]. Models [109]. Modules [117]. Monads [81]. Moped [23]. Multi [123]. Multi-Level [123]. Multilisp [89]. Multimethods [39]. Mystery [1].

Names [109]. Necessary [18]. need [82]. Next [75]. Non [1]. Non-Reflective [1]. Nucleic [89].

Oaklisp [2]. Object [2, 9–12, 17, 19, 45, 46, 108]. Object-Based [108]. Object-Oriented [2, 9, 19, 45, 46]. Objects [43]. OOP [106]. Operational [120, 128]. Optimizations [7]. Optimizing [46]. OR-Parallel [88]. Order [115]. Organizing [44]. Oriented [2, 9, 19, 45, 46]. Overloading [100]. Overview [10, 66].

P [118]. pages [4, 15, 25, 33]. Parallel [37, 52–54, 56, 63, 84, 87, 88]. Parallellization [20]. Parents [43]. Partial [13, 63, 99, 100, 113, 122, 143]. Parts [43]. Passing [3, 59, 77, 82]. Performance [28]. PERs [101]. Persistence [119]. Persistent [118]. Phylogenetic [88]. Plurals [72]. Polymorphic [78, 142]. Polymorphism [35, 105, 120]. Portable [23]. power [42]. Practical [70]. Predicate [58]. Prentice [33]. Preprocessor [29]. PreScheme [96].

Problems [89]. Procedure [58]. Procedures [30]. Processing [52, 54, 56]. Program [36, 123]. Programmer [11, 12, 19]. Programming [4, 9, 14–16, 19, 31, 56, 57, 117]. Programs [20, 22, 28, 32, 37, 44, 46, 53, 77, 84]. Projections [101, 102]. Prolog [88]. Protocol [68, 119]. Prototype [111]. Prototype-Based [111]. Prototypes [45]. Prototyping [8]. Publishing [25, 41]. Pure [16]. Purpose [52].

Rapid [8]. Reasoning [77]. Recursion [80]. Reference [25, 55, 124]. Reference-Counting [124]. Reflection [114]. Reflective [1, 111–113]. Registers [59]. Relative [22]. Reliable [117]. Remarks [50]. Removal [115]. Research [124]. Reuse [22]. Revealed [1]. Review [4, 15, 19, 25, 33, 41]. Revisited [40]. Rewriting [145]. Richard [4]. Riesbeck [15]. Runtime [18, 34].

S [41]. Satisfaction [87, 89]. save [59]. Scalable [52]. Scan [27]. Schemata [76]. Scheme [2, 20, 26, 31, 54, 55, 62, 91, 92, 94, 95, 126, 144]. Self [13, 42, 43, 45]. Self-Applicable [13]. Semantics [108, 111, 120, 128]. Separation [5]. Several [28]. Shared [43]. SIMD [72]. Simple [105, 112]. simplicity [42]. Softcover [25, 33]. Solving [89]. Sonya [19]. Special [107]. Specialization [123]. Specification [11, 12, 17, 22]. Splitting [46]. Standard [35, 98]. State [104, 106, 108]. Static [102]. Storage [116]. Strictness [110]. Structure [89]. Structures [85, 91]. Style [3, 59, 77, 82]. Subcontinuations [83]. SUBTYPEP [58]. Support [39, 119]. Symbolic [41]. Syntactic [62]. System [10–12, 17, 34, 54, 87, 88, 90, 95]. Systems [28].

Tags [18]. Technical [5]. techniques [145].

Telos [69]. Their [24]. Theory [102, 125]. Threads [71, 127]. Time [98, 101, 114]. Time-Dependent [114]. Title [141]. Toolbox [71]. Touretzky [41]. Tower [1]. Trace [128]. Trace-Based [128]. Tractable [92]. Transformational [22]. Trivial [125]. TS [91]. TS/Scheme [91]. Tutor [8]. Tutorial [33]. Two [28]. Type [35, 46, 70, 78, 102, 142]. Typed [45, 46, 106].

UCL [118]. Unboxed [120]. understanding [113]. Unnecessary [108]. Using [89].

Value [5]. Variable [30]. Variables [32, 38]. Verified [94–96, 144]. Version [49, 50]. via [119]. Visualizing [126]. VLISP [94–96, 107, 144]. Volume [130–139].

Wendy [33]. Wesley [19, 25]. Without [44].

Yield [127].

References

Wand:1988:MTR

[1] Mitchell Wand and Daniel P. Friedman. The mystery of the tower revealed: A non-reflective description of the reflective tower. Lisp and Symbolic Computation, 1(1):11–37, June 1988. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). Reprinted in Meta-Level Architectures and Reflection (P. Maes and D. Nardi, eds.) North-Holland, Amsterdam, 1988, pp. 111– 134. Preliminary version appeared in Proc. 1986 ACM Conf. on Lisp and Functional Programming, 298–307.

Lang:1988:OOO

[2] Kevin J. Lang and Barak A. Pearlmutter. Oaklisp: An object-oriented dialect of Scheme. Lisp and Symbolic Computation, 1(1):39–51, June 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Dybvig:1988:EPS

[3] R. Kent Dybvig, Daniel P. Friedman, and Christopher T. Haynes. Expansionpassing style: A general macro mechanism. Lisp and Symbolic Computation, 1(1):53–75, June 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1988:BRL

[4] Anonymous. Book review: Lisp Lore: A Guide to Programming the Lisp Machine, 2d ed., Hank Bromley and Richard Lamson, Boston: Kluwer. 1987. 337 pages hardbound. $47.50. Lisp and Symbolic Computation, 1(1):77–79, June 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Gabriel:1988:ETI

[5] Richard P. Gabriel and Kent M. Pitman. Endpaper: Technical issues of separation in function cells and value cells. Lisp and Symbolic Computation, 1(1): 81–101, June 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

ODonnell:1988:DAL

[6] John T. O’Donnell and Cordelia V. Hall. Debugging in applicative languages. Lisp and Symbolic Computation, 1(2):113–145, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bloss:1988:COL

[7] Adrienne Bloss, Paul R. Hudak, and Jonathan H. Young. Code optimizations for lazy evaluation. Lisp and Symbolic Computation, 1(2):147–164, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Giuse:1988:LRP

[8] Dario Giuse. LISP as a rapid prototyping environment: The Chinese tutor. Lisp and Symbolic Computation, 1 (2):165–184, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Lieberherr:1988:OOP

[9] Karl J. Lieberherr. Object-oriented programming with class dictionaries. Lisp and Symbolic Computation, 1(2): 185–212, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Demichiel:1988:OCL

[10] Linda G. Demichiel. Overview: The Common Lisp Object System. Lisp and Symbolic Computation, 1(2):227–244, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bobrow:1988:CLOa

[11] Daniel G. Bobrow, Linda G. Demichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common Lisp Object System specification 1. programmer interface concepts. Lisp and Symbolic Computation, 1(2): 245–298, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bobrow:1988:CLOb

[12] Daniel G. Bobrow, Linda G. Demichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common Lisp Object System specification 2. functions in the programmer interface. Lisp and Symbolic Computation, 1(2):299–394, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Jones:1989:MSA

[13] Neil D. Jones, Peter Sestoft, and Harald Søndergaard. Mix: A self-applicable partial evaluator for experiments in compiler generation. Lisp and Symbolic Computation, 2(1):9–50, February 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). DIKU Report 91/12.

Ait-kaci:1989:ILF

[14] Hassan Ait-kaci and Roger Nasr. Integrating logic and functional programming. Lisp and Symbolic Computation, 2(1):51–89, February 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1989:BRA

[15] Anonymous. Book review: Artificial Intelligence Programming, 2d ed. Charniak, Riesbeck, McDermott, and Meehan. New Jersey: Lawrence Erlbaum Associates, 1987. 533 pages hardbound. Lisp and Symbolic Computation, 2(1):91–93, February 1989. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Grue:1989:APF

[16] Klaus E. Grue. Arrays in pure functional programming languages. Lisp and Symbolic Computation, 2(1):105– 113, February 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Olthoff:1989:ASM

[17] Walter Olthoff and James Kempf. An algebraic specification of method combination for the Common Lisp Object System. Lisp and Symbolic Computation, 2(1):115–152, February 1989. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Appel:1989:RTA

[18] Andrew W. Appel. Runtime tags aren’t necessary. Lisp and Symbolic Computation, 2(2):153–162, June 1989. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). I have the techreport CS-TR-142-88, Princeton University, March 88.

Anonymous:1989:BRO

[19] Anonymous. Book review: Object-Oriented Programming in Common LISP: A Programmer’s Guide to CLOS, Sonya E. Keene, Massachusetts: Addison-Wesley, 1989. Lisp and Symbolic Computation, 2(2):163–165, June 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Harrison:1989:IAA

[20] Williams Ludwell Harrison, III. The interprocedural analysis and automatic parallellization of Scheme programs. Lisp and Symbolic Computation, 2(3/4): 179–396, October 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Zorn:1990:DFC

[21] Benjamin Zorn and Paul Hilfinger. Direct function calls in Lisp. Lisp and Symbolic Computation, 3(1):13–20, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Runciman:1990:RST

[22] Colin Runciman and Nigel Jagger. Relative specification and transformational reuse of functional programs. Lisp and Symbolic Computation, 3(1):21– 37, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Pourheidari:1990:MPD

[23] Mohammad Pourheidari, Robert R. Kessler, and Harold Carr. Moped (A portable debugger). Lisp and Symbolic Computation, 3(1):39–65, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Sitaram:1990:CDT

[24] Dorai Sitaram and Matthias Felleisen. Control delimiters and their hierarchies. Lisp and Symbolic Computation, 3(1):67–99, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Touretzky:1990:BRC

[25] David S. Touretzky. Book review: Common Lisp: the Reference, by Franz, Inc. Addison-Wesley Publishing Company, 1988. softcover; 919 pages. Lisp and Symbolic Computation, 3(1):101– 102, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bonzon:1990:MEL

[26] Pierre E. Bonzon. A metacircular evaluator for a logical extension of Scheme. Lisp and Symbolic Computation, 3(2): 113–133, April 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Derbyshire:1990:MSG

[27] Margaret H. Derbyshire. Mark scan garbage collection on a distributed architecture. Lisp and Symbolic Computation, 3(2):135–170, April 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Acuff:1990:EPT

[28] Richard Acuff. Endpaper: Performance of two Common Lisp programs on several systems. Lisp and Symbolic Computation, 3(2):171–204, April 1990. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Diamant:1990:PBI

[29] John R. Diamant and Gerald P. Duggan. A preprocessor based implementation of Common Lisp. Lisp and Symbolic Computation, 3(3):211–228, September 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Dybvig:1990:NAP

[30] R. Kent Dybvig and Robert Hieb. A new approach to procedures with variable arity. Lisp and Symbolic Computation, 3 (3):229–244, September 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Ruf:1990:LIL

[31] Erik Ruf and Daniel Weise. LogScheme: Integrating logic programming into Scheme. Lisp and Symbolic Computation, 3(3):245–288, September 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bage:1990:CEF

[32] Göran Bage and Gary Lindstrom. Combinator evaluation of functional programs with logical variables. Lisp and Symbolic Computation, 3(3):289–320, September 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1990:BRC

[33] Anonymous. Book review: Common Lisp: A Tutorial, by Wendy L. Milner. Prentice Hall, 1988. softcover; 519 pages. Lisp and Symbolic Computation, 3(3):321–325, September 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Appel:1990:RS

[34] Andrew W. Appel. A runtime system. Lisp and Symbolic Computation, 3(4):343–380, November 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Ghosh-Roy:1990:CSM

[35] R. Ghosh-Roy. Conjunction-type Standard ML polymorphism. Lisp and Symbolic Computation, 3(4):381–409, November 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kessler:1991:LAP

[36] Robert R. Kessler and Jed J. Krohnfeldt. A logical approach to program analysis. Lisp and Symbolic Computation, 4(1):5–27, January 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Larus:1991:CLP

[37] James R. Larus. Compiling Lisp programs for parallel execution. Lisp and Symbolic Computation, 4(1):29– 99, January 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Miller:1991:FVF

[38] James S. Miller and Guillermo J. Rozas. Free variables and first-class environments. Lisp and Symbolic Computation, 4(2):107–141, April 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Koschmann:1991:DBS

[39] Timothy Koschmann. Designing a browser to support multimethods and method combination. Lisp and Symbolic Computation, 4(2):143–154, April 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Fateman:1991:EFB

[40] Richard J. Fateman. Endpaper: FR-POLY: A benchmark revisited. Lisp and Symbolic Computation, 4(2):155– 164, April 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

White:1991:BRK

[41] Jon L. White. Book review: A “Kinder, Gentler LISP?” COMMON LISP: A Gentle Introduction to Symbolic Computation, by David S. Touretzky. Benjamin/Cummings Publishing Company 1990. Lisp and Symbolic Computation, 4(2):165–177, April 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Ungar:1991:SPS

[42] David Ungar and Randall B. Smith. SELF. the power of simplicity. Lisp and Symbolic Computation, 4(3):187– 205, July 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). Preliminary version appeared in Proc. ACM Symp. on Object-Oriented Programming: Systems, Languages, and Applications, 1987, 227-241.

Chambers:1991:PSP

[43] Craig Chambers, David Ungar, Bay-Wei Chang, and Urs Hölzle. Parents are shared parts of objects: Inheritance and encapsulation in SELF. Lisp and Symbolic Computation, 4(3):207–222, July 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Ungar:1991:OPC

[44] David Ungar, Craig Chambers, Bay-Wei Chang, and Urs Hölzle. Organizing programs without classes. Lisp and Symbolic Computation, 4(3):223–242, July 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Chambers:1991:EIS

[45] Craig Chambers, David Ungar, and Elgin Lee. An efficient implementation of SELF, a dynamically-typed objectoriented language based on prototypes. Lisp and Symbolic Computation, 4(3): 243–281, July 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Chambers:1991:ITA

[46] Craig Chambers and David Ungar. Iterative type analysis and extended message splitting: Optimizing dynamicallytyped object-oriented programs. Lisp and Symbolic Computation, 4(3):283– 310, July 1991. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1992:Ia

[47] Anonymous.  Introduction.  Lisp

and Symbolic Computation, 4(4):317– ??, March 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Cengarle:1992:ADK

[48] María Victoria Cengarle, Luis Mandel, Martin Wirsing, Heiner Brand, Klaus Däßler, and Thekla Schneider. An approach to the DIN Kernel Lisp definition. Lisp and Symbolic Computation, 4(4):319–369, March 1992. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Baker:1992:CDK

[49] Henry G. Baker. Critique of DIN Kernel Lisp definition version 1.2. Lisp and Symbolic Computation, 4(4):371– 398, March 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Cengarle:1992:ABR

[50] María Victoria Cengarle, Luis Mandel, Martin Wirsing, Heiner Brand, Klaus Däßler, and Thekla Schneider. Answer to Baker’s remarks of July 1991 on version 1.0. Lisp and Symbolic Computation, 4(4):399–404, March 1992. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1992:Ib

[51] Anonymous.  Introduction.  Lisp

and Symbolic Computation, 5(1/2):5– ??, May 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Davis:1992:MGP

[52] Al Davis. Mayfly: A general-purpose, scalable, parallel processing architecture. Lisp and Symbolic Computation, 5(1/2):7–48, May 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Conery:1992:PLP

[53] John S. Conery. Parallel logic programs on the Mayfly. Lisp and Symbolic Computation, 5(1/2):49–72, May 1992. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kessler:1992:ICS

[54] R. Kessler, H. Carr, L. Stoller, and M. Swanson. Implementing Concurrent Scheme for the Mayfly distributed parallel processing system. Lisp and Symbolic Computation, 5(1/2):73–93, May 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Swanson:1992:CSR

[55] Mark R. Swanson. Concurrent Scheme reference. Lisp and Symbolic Computation, 5(1/2):95–104, May 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Evans:1992:DMP

[56] John D. Evans and Robert R. Kessler. DPOS: A metalanguage and programming environment for parallel processing. Lisp and Symbolic Computation, 5(1/2):105–123, May 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Gang:1992:FLP

[57] Cheng Gang and Yun-Zheng Zhang. A “functional + logic” programming language in interpretation-compilation implementation. Lisp and Symbolic Computation, 5(3):133–156, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Baker:1992:DPC

[58] Henry G. Baker. A decision procedure for Common Lisp’s SUBTYPEP predicate. Lisp and Symbolic Computation, 5 (3):157–190, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Appel:1992:CSR

[59] Andrew W. Appel and Zhong Shao. Callee-save registers in continuationpassing style. Lisp and Symbolic Computation, 5(3):191–221, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

URL ftp://ftp.cs.princeton.edu/
reports/1991/326.ps.Z.

Borning:1992:CH

[60] Alan Borning, Bjorn N. Freeman-Benson, and Molly Wilson. Constraint hierarchies. Lisp and Symbolic Computation, 5(3):223–270, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Lapalme:1992:DEL

[61] Guy Lapalme and Mario Latendresse. A debugging environment for lazy functional languages. Lisp and Symbolic Computation, 5(3):271–287, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Dybvig:1992:SAS

[62] R. Kent Dybvig, Robert Hieb, and Carl Bruggeman. Syntactic abstraction in Scheme. Lisp and Symbolic Computation, 5(4):295–326, December 1992. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Consel:1992:PEP

[63] Charles Consel and Olivier Danvy. Partial evaluation in parallel. Lisp and Symbolic Computation, 5(4):327–342, December 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Lee:1992:FE

[64] Shinn-Der Lee and Daniel P. Friedman. First-class extents. Lisp and Symbolic Computation, 5(4):343–375, December 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1993:E

[65] Anonymous. Editorial. Lisp and Symbolic Computation, 6(1/2):5–??, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Padget:1993:OE

[66] Julian Padget, Greg Nuyens, and Harry Bretthauer. An overview of EULISP. Lisp and Symbolic Computation, 6(1/2): 9–98, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bradford:1993:EE

[67] Russell J. Bradford and David C. De Roure. EuLisp in education. Lisp and Symbolic Computation, 6(1/2):99– 118, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bretthauer:1993:BEM

[68] Harry Bretthauer, Jürgen Kopp, Harley E. Davis, and Keith J. Playford. Balancing the EuLisp metaobject protocol. Lisp and Symbolic Computation, 6(1/2):119–138, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Broadbery:1993:AT

[69] Peter A. Broadbery and Christopher Burdorf. Applications of Telos. Lisp and Symbolic Computation, 6(1/2):139– 158, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kind:1993:PAT

[70] Andreas Kind and Horst Friedrich. A practical approach to type inference for EuLisp. Lisp and Symbolic Computation, 6(1/2):159–176, August 1993. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Berrington:1993:ETC

[71] Neil Berrington, Peter A. Broadbery, David De Roure, and Julian A. Padget. EuLisp threads: A concurrency toolbox. Lisp and Symbolic Computation, 6(1/2):177–200, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL ftp://ftp.

                                        bath.ac.uk/pub/eulisp/threads.ps.
                                        gz.

Merrall:1993:PSE

[72] Simon Merrall and Julian A. Padget. Plurals: A SIMD extension to Eu-Lisp. Lisp and Symbolic Computation, 6(1/2):201–220, August 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1993:I

[73] Anonymous. Introduction. Lisp and Symbolic Computation, 6(3/4):229–??, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Reynolds:1993:DC

[74] John C. Reynolds. The discoveries of continuations. Lisp and Symbolic Computation, 6(3/4):233–248, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Morris:1993:NFL

[75] F. Lockwood Morris. The next 700 formal language descriptions. Lisp and Symbolic Computation, 6(3/4):249–258, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). Original manuscript dated November 1970.

Fischer:1993:LCS

[76] Michael J. Fischer. Lambda-calculus schemata. Lisp and Symbolic Computation, 6(3/4):259–288, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Sabry:1993:RAP

[77] Amr Sabry and Matthias Felleisen. Reasoning about programs in continuationpassing style. Lisp and Symbolic Computation, 6(3/4):289–360, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Harper:1993:PTA

[78] Robert W. Harper and Mark Lillibridge. Polymorphic type assignment and CPS conversion. Lisp and Symbolic Computation, 6(3/4):361–380, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). See corrigendum [142].

Anonymous:1994:I

[79] Anonymous.  Introduction.  Lisp

and Symbolic Computation, 7(1):7–??, January 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Filinski:1994:RI

[80] Andrzej Filinski. Recursion from iteration. Lisp and Symbolic Computation, 7(1):11–38, January 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Wadler:1994:MCC

[81] Philip Wadler. Monads and composable continuations. Lisp and Symbolic Computation, 7(1):39–56, January 1994. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Okasaki:1994:CNC

[82] Chris Okasaki, Peter Lee, and David Tarditi. Call-by-need and continuationpassing style. Lisp and Symbolic Computation, 7(1):57–82, January 1994. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL ftp:

//ftp.cs.cmu.edu/usr/anon/user/
cokasaki/lasc94.dvi.gz.

Hieb:1994:S

[83] Robert Hieb, Kent Dybvig, and Claude W. Anderson, III. Subcontinuations. Lisp and Symbolic Computation, 7(1):83–110, January 1994. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Conery:1994:CBC

[84] John S. Conery. Continuation based control in the implementation of parallel logic programs. Lisp and Symbolic Computation, 7(1):111–134, January 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Chakrabarti:1994:DDS

[85] Soumen Chakrabarti and Katherine Yelick. Distributed data structures and algorithms for Gröbner basis computation. Lisp and Symbolic Computation, 7(2/3):147–172, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Weeks:1994:CAI

[86] Stephen Weeks, Suresh Jagannathan, and James Philbin. A concurrent abstract interpreter. Lisp and Symbolic Computation, 7(2/3):173–193, ????

1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Ho:1994:CPC

[87] Kinson Ho, Hans W. Guesgen, and Paul N. Hilfinger. CONSAT: A parallel constraint satisfaction system. Lisp and Symbolic Computation, 7(2/3):195– 210, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Matsuda:1994:AOP

[88] Hideo Matsuda and Yukio Kaneda. An application of an OR-parallel Prolog system to phylogenetic analysis. Lisp and Symbolic Computation, 7(2/3):211– 229, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Feeley:1994:UMS

[89] Marc Feeley, Marcel Turcotte, and Guy Lapalme. Using Multilisp for solving constraint satisfaction problems: an application to nucleic acid 3D structure determination. Lisp and Symbolic Computation, 7(2/3):231–247, ???? 1994. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Karp:1994:GCG

[90] Peter D. Karp, John D. Lowrance, Thomas M. Strat, and David E. Wilkins. The Grasper-CL graph management system. Lisp and Symbolic Computation, 7(4):251–290, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Jagannathan:1994:TSD

[91] Suresh Jagannathan. TS/Scheme: distributed data structures in Lisp. Lisp and Symbolic Computation, 7(4):291– 314, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kelsey:1994:TSI

[92] Richard A. Kelsey and Jonathan A. Rees. A tractable Scheme implementation. Lisp and Symbolic Computation, 7(4):315–335, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Appel:1994:LHC

[93] Andrew W. Appel. Loop headers in λ-calculus or CPS. Lisp and Symbolic Computation, 7(4):337–343, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Guttman:1995:VVIa

[94] Joshua D. Guttman, John D. Ramsdell, and Mitchell Wand. VLISP: A verified implementation of Scheme. Lisp and Symbolic Computation, 8 (1/2):5–32, March 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL ftp:/

/ftp.ccs.neu.edu/pub/people/wand/
vlisp/lasc/overview.dvi;   http:/
/www.wkap.nl/oasis.htm/81420.

Guttman:1995:VVS

[95] Joshua D. Guttman, John D. Ramsdell, and Vipin Swarup. The VLISP verified Scheme system. Lisp and Symbolic Computation, 8(1/2):33–110, March 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/81421.

Oliva:1995:VVP

[96] Dino P. Oliva, John D. Ramsdell, and Mitchell Wand. The VLISP verified PreScheme compiler. Lisp and Symbolic Computation, 8(1/2):111–182, March 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

URL ftp://ftp.ccs.neu.edu/pub/
people/wand/vlisp/lasc/prescheme.
dvi;   http://www.wkap.nl/oasis.
htm/81422.

Sestoft:1995:GEF

[97] Peter Sestoft and Harald Søndergaard. Guest editors’ foreword. Lisp and Symbolic Computation, 8(3):189–??, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Birkedal:1995:BTA

[98] Lars Birkedal and Morten Welinder. Binding-time analysis for Standard ML. Lisp and Symbolic Computation, 8(3): 191–208, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

wkap.nl/oasis.htm/92449.

Danvy:1995:EEE

[99] Olivier Danvy, Karoline Malmkjær, and Jens Palsberg. The essence of eta-expansion in partial evaluation. Lisp and Symbolic Computation, 8(3): 209–227, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

daimi.aau.dk/~danvy/Papers/danvy-
malmkjaer-palsberg-lasc95.html;
ftp://ftp.daimi.aau.dk/pub/danvy/
Papers/danvy-malmkjaer-palsberg-
lasc95.ps.gz; http://www.wkap.nl/
oasis.htm/92450.

Jones:1995:DFO

[100] Mark P. Jones. Dictionary-free overloading by partial evaluation. Lisp and Symbolic Computation, 8(3):229–248, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/92451.

Davis:1995:PPB

[101] Kei Davis. PERs from projections for binding time analysis. Lisp and Symbolic Computation, 8(3):249–266, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/92452.

Baker-Finch:1995:TTP

[102] Clement A. Baker-Finch. Type theory and projections for static analysis. Lisp and Symbolic Computation, 8 (3):267–285, September 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

                                        wkap.nl/oasis.htm/92453.

Mason:1995:I

[103] Ian A. Mason and Martin Odersky. Introduction. Lisp and Symbolic Computation, 8(4):291–292, December 1995. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

                                        www.wkap.nl/oasis.htm/97402.

Launchbury:1995:SH

[104] John Launchbury and Simon L. Peyton Jones. State in Haskell. Lisp and Symbolic Computation, 8(4):293– 341, December 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/97403.

Wright:1995:SIP

[105] Andrew K. Wright. Simple imperative polymorphism. Lisp and Symbolic Computation, 8(4):343–355, December 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/97404.

Eifrig:1995:ITO

[106] Jonathan Eifrig, Scott Smith, Valery Trifonov, and Amy Zwarico. An interpretation of typed OOP in a language with state. Lisp and Symbolic Computation, 8(4):357–397, December 1995. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

www.wkap.nl/oasis.htm/97405.

Guttman:1995:SIV

[107] Joshua D. Guttman and Mitchell Wand.

Special issue on VLISP. Lisp and Symbolic Computation, 8(1/2):??, March 1995. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Reddy:1996:GSC

[108] Uday S. Reddy. Global state considered unnecessary: An introduction to object-based semantics. Lisp and Symbolic Computation, 9(1):7–76, February 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/105832.

Stark:1996:CML

[109] Ian Stark. Categorical models for local names. Lisp and Symbolic Computation, 9(1):77–107, February 1996. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

www.wkap.nl/oasis.htm/105834.

Springer:1996:SAI

[110] Jonathan Springer and Samuel N. Kamin. Strictness analysis in the imperative lambda calculus. Lisp and Symbolic Computation, 9(1):109–143, February 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/105837.

Malenfant:1996:SIR

[111] Jacques Malenfant, Christophe Dony, and Pierre Cointe. A semantics of introspection in a reflective prototype-based language. Lisp and Symbolic Computation, 9(2/3):153–179, May/June 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/111830.

Jefferson:1996:SRI

[112] Stanley Jefferson and Daniel P. Friedman. A simple reflective interpreter. Lisp and Symbolic Computation, 9(2/3): 181–202, May/June 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

                                        wkap.nl/oasis.htm/111833.

Asai:1996:DPE

[113] Kenichi Asai, Satoshi Matsuoka, and Akinori Yonezawa. Duplication and partial evaluation — for a better understanding of reflective languages -. Lisp and Symbolic Computation, 9(2/3): 203–241, May/June 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

                                        wkap.nl/oasis.htm/111834.

Honda:1996:RTD

[114] Yasuaki Honda and Mario Tokoro. Reflection and time-dependent computing: Experiences with the R^{2} architecture. Lisp and Symbolic Computation, 9(2/3): 243–277, May/June 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

wkap.nl/oasis.htm/111837.

Chin:1996:HOR

[115] Wei-Ngan Chin and John Darlington.

A higher-order removal method. Lisp and Symbolic Computation, 9(4):287– 322, December 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/119671.

Tarditi:1996:MCS

[116] David Tarditi and Amer Diwan. Measuring the cost of storage management. Lisp and Symbolic Computation, 9(4):323– 342, December 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/119672.

Tung:1996:RIP

[117] Sho-Huan Simon Tung and R. Kent Dybvig. Reliable interactive programming with modules. Lisp and Symbolic Computation, 9(4):343–358, December 1996. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/119673.

Jacobs:1997:UDI

[118] J. H. Jacobs and M. R. Swanson.

UCL+P — defining and implementing persistent Common Lisp. Lisp and Symbolic Computation, 10(1):5– 38, June 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/134677.

Lee:1997:ASP

[119] Arthur H. Lee and Joseph L. Zachary. Adding support for persistence to CLOS via its metaobject protocol. Lisp and Symbolic Computation, 10(1):39– 60, June 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/134678.

Ohori:1997:UOS

[120] Atsushi Ohori and Tomonobu Takamizawa. An unboxed operational semantics for ML polymorphism. Lisp and Symbolic Computation, 10(1):61–91, June 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/134679.

Kessler:1997:E

[121] Bob Kessler and Carolyn Talcott. Editorial. Lisp and Symbolic Computation, 10(2):99, July 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/142927.

Danvy:1997:PEE

[122] Olivier Danvy and Mayer Goldberg. Partial evaluation of the Euclidian algorithm. Lisp and Symbolic Computation, 10(2):101–111, July 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

wkap.nl/oasis.htm/136321.

Gluck:1997:APG

[123] Robert Glück and Jesper Jørgensen.

An automatic program generator for multi-level specialization. Lisp and Symbolic Computation, 10(2):113–158, July 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/136322.

Wise:1997:RDH

[124] David S. Wise, Brian Heck, Caleb Hess, Willie Hunt, and Eric Ost. Research demonstration of a hardware referencecounting heap. Lisp and Symbolic Computation, 10(2):159–181, July 1997. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

www.wkap.nl/oasis.htm/136324.

Wand:1998:TFT

[125] Mitchell Wand. The theory of Fexprs is trivial. Lisp and Symbolic Computation, 10(3):189–199, May 1998. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

www.wkap.nl/oasis.htm/161782.

Tung:1998:VES

[126] Sho-Huan Simon Tung. Visualizing evaluation in Scheme. Lisp and Symbolic Computation, 10(3):201–222, May 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/161783.

Kumar:1998:TYC

[127] Sanjeev Kumar, Carl Bruggeman, and R. Kent Dybvig. Threads yield continuations. Lisp and Symbolic Computation, 10(3):223–236, May 1998. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://

                                        www.wkap.nl/oasis.htm/161784.

Schmidt:1998:TBA

[128] David A. Schmidt. Trace-based abstract interpretation of operational semantics. Lisp and Symbolic Computation, 10(3):237–271, May 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

                                        wkap.nl/oasis.htm/161785.

Kessler:1998:E

[129] Bob Kessler and Carolyn Talcott. Editorial. Lisp and Symbolic Computation, 10(4):279, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). URL http://www.

                                        wkap.nl/oasis.htm/167760.

Anonymous:1998:Va

[130] Anonymous. Volume 1 (1988). Lisp and Symbolic Computation, 10(4):281– 287, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/167761.

Anonymous:1998:Vb

[131] Anonymous. Volume 2 (1989). Lisp and Symbolic Computation, 10(4):289– 293, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/167762.

Anonymous:1998:Vc

[132] Anonymous. Volume 3 (1990). Lisp and Symbolic Computation, 10(4):295– 303, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

                                        tronic). URL http://www.wkap.nl/
                                        oasis.htm/167763.

Anonymous:1998:Vd

[133] Anonymous. Volume 4 (1991). Lisp and Symbolic Computation, 10(4):305– 313, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167764.

Anonymous:1998:Ve

[134] Anonymous. Volume 5 (1992). Lisp and Symbolic Computation, 10(4):315– 323, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167766.

Anonymous:1998:Vf

[135] Anonymous. Volume 6 (1993). Lisp and Symbolic Computation, 10(4):325– 332, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167767.

Anonymous:1998:Vg

[136] Anonymous. Volume 7 (1994). Lisp and Symbolic Computation, 10(4):333– 342, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167768.

Anonymous:1998:Vh

[137] Anonymous. Volume 8 (1995). Lisp and Symbolic Computation, 10(4):343– 349, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167769.

Anonymous:1998:Vi

[138] Anonymous. Volume 9 (1996). Lisp and Symbolic Computation, 10(4):351– 357, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167770.

Anonymous:1998:Vj

[139] Anonymous. Volume 10 (1997). Lisp and Symbolic Computation, 10(4):359– 366, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/167771.

Anonymous:1998:AI

[140] Anonymous. Author index. Lisp and Symbolic Computation, 10(4):367– 369, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/186923.

Anonymous:1998:TI

[141] Anonymous. Title index. Lisp and Symbolic Computation, 10(4):371–373, August 1998. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.wkap.nl/
oasis.htm/186924.

Harper:2003:CPT

[142] Robert Harper and Mark Lillibridge. Corrigendum: Polymorphic type assignment and CPS conversion. Higher-Order and Symbolic Computation, 16(4):401, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/24/5/abstract.htm; http://
ipsapp008.kluweronline.com/content/
getfile/4979/24/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&

volume=16&issue=4&spage=401. See

[78].

Jones:1985:EPE

[143] N. D. Jones, P. Sestoft, and H. Søndergaard.

An experiment in partial evaluation: The generation of a compiler generator. In Jouannaud [145], pages 124–140. CODEN LNCSD9. ISBN 0-387-15976-2 (paperback: U.S.). ISSN 0302-9743. LCCN QA76.6 .R4641 1985. Papers presented at the First International Conference on Rewriting Techniques and Applications, held in Dijon, Burgundy, France, May 20–22, 1985.

Guttman:1995:VVIb

[144] Joshua D. Guttman and Mitchell Wand, editors. VLISP: A Verified Implementation of Scheme. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-9566-2. 184 pp. LCCN QA76.73.L23 V55

1995. URL ftp://ftp.ccs.neu.edu/
pub/people/wand/vlisp/lasc. Origi-

nally published as a special double issue of the journal Lisp and Symbolic Computation (Volume 8, Issue 1/2).

Jouannaud:1985:RTA

[145] Jean-Pierre Jouannaud, editor. Rewriting techniques and applications: Dijon, France, May 20–22, 1985, volume 202 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNCSD9. ISBN 0-387-15976-2 (paperback: U.S.). ISSN 0302-9743. LCCN QA76.6 .R4641 1985. Papers presented at the First International Conference on Rewriting Techniques and Applications, held in Dijon, Burgundy, France, May 20–22, 1985.